I need help on steps for solving a system of equations in substituting,elimination ,and graphing method for problem:
X+2y=10 3x-y=9
Thank you,
I tried
2y=x+10 y=-1\2 x10
3x-(1\5x15)=9
3x+1\3x-5=9
3.5x-5=9
(3,14)

Answer by Alan3354(69443)

X+2y=10 3x-y=9
Thank you,
I tried
2y=x+10 ****** 2y = -x + 10
y=-1\2 x10 ***** Not sure what this means, but y = -x/2 + 5
3x-(1\5x15)=9
3x+1\3x-5=9
3.5x-5=9
(3,14)
===============
X+2y=10
3x-y=9
Don't make any fractions.
3x-y=9 --> y = 3x - 9
Sub for y in the 1st eqn
--
x + 2y = 10
x + 2(3x-9) = 10
x + 6x-18 = 10
7x = 28
x = 4